OpenStudy (anonymous):

Solve the equation by completing the square. x^2+12x+18=0. Please explain what completing the square means.

OpenStudy (asnaseer):

do you know how to solve using quadractic formulae (not that this is how you should do this - I just want to know if you are aware of it)?

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

ok, so the quadratic formulae can be derived by whats called "completing the square"

OpenStudy (asnaseer):

in general, we have:\[ax^2+bx+c=0\]we first divide by 'a' to get:\[x^2+\frac{bx}{a}+\frac{c}{a}=0\]

OpenStudy (asnaseer):

the next step is to just concentrate on the x^2 and x terms - forget the constant

OpenStudy (asnaseer):

and we see if this can be written in the form (x+q)^2

OpenStudy (asnaseer):

this is what "completing the square" is about

OpenStudy (anonymous):

how?

OpenStudy (asnaseer):

let me show you....

OpenStudy (anonymous):

6

OpenStudy (anonymous):

36

OpenStudy (asnaseer):

\[(x+\frac{b}{2a})^2=x^2+\frac{bx}{a}+\frac{b^2}{4a^2}\]

OpenStudy (asnaseer):

@tapdancer8: do you understand the expansion I just wrote?

OpenStudy (anonymous):

6^2 = 36 good, now add 36 to the right side like this x^2+12x + 36 +18 Now can you factor the x^2+12x + 36 part?

OpenStudy (anonymous):

@asnaseer: sort of, it's an expanded version of the quadratic formula

OpenStudy (asnaseer):

we had this earlier:\[x^2+\frac{bx}{a}+\frac{c}{a}=0\]and I just showed you that:\[(x+\frac{b}{2a})^2=x^2+\frac{bx}{a}+\frac{b^2}{4a^2}\]so we can combine the two next

OpenStudy (anonymous):

@moses_:yes, it becomes (x+6)^2

OpenStudy (asnaseer):

\[(x+\frac{b}{2a})^2-\frac{b^2}{4a^2}=x^2+\frac{bx}{a}\]so put this into our 1st equation to get:\[(x+\frac{b}{2a})^2-\frac{b^2}{4a^2}+\frac{c}{a}=0\]

OpenStudy (asnaseer):

then rewrite as:\[(x+\frac{b}{2a})^2=\frac{b^2}{4a^2}-\frac{c}{a}\]

OpenStudy (asnaseer):

this is now in the form of: (something)^2 = constant and your next step would be: something = squareroot(constant)

OpenStudy (anonymous):

good, thats essentially what asnaseer is trying to show you, but you can see it faster this way. The whole point is so you can now factor a part of the left hand side of the equation. Now you also must subtract 36 on the left hand side of the equation so it comes out to be the exact same as when we started so you have x^2+12x + 36 +18-36. No you already factored the left hand side first 3 terms and as you can see by subtracting the 36 you have not really done anything illegal in terms of the rules of math. so you turned x^2+12x + 36 into (x+6)^2 and you are left with 18-36 so what is that all combined? in the new form?

OpenStudy (asnaseer):

this is what completing the square is about. so, in your case:\[x^2+12x+18=0\]you would first see that:\[(x+6)^2=x^2+12x+36\]and use this to get:\[(x+6)^2-18=0\]which leads to:\[(x+6)^2=18\]and solve from there

OpenStudy (asnaseer):

so completing the square essentially involves: 1) divide by the coefficient of x^2 2) take half the resulting coefficient of x to form the (x+p)^2 term

OpenStudy (asnaseer):

in your case the coefficient of x^2 was just 1 - so no need to do any initial division

OpenStudy (anonymous):

\[\left( x+6 \right)^2=-18\]\[\sqrt{\left( x+6 \right)^2}=\sqrt{-18}\]\[x+6=\pm i \sqrt{18}\]\[x=-6\pm 3i \sqrt{2}\]

OpenStudy (anonymous):

right?

OpenStudy (asnaseer):

you have the sign wrong

OpenStudy (anonymous):

which one?

OpenStudy (asnaseer):

\[(x+6)^2=18\]

OpenStudy (anonymous):

the -18 is 18

OpenStudy (anonymous):

but you just solved his equation...you should really take a look at how he got there.

OpenStudy (anonymous):

how? if I had 18-36?

OpenStudy (asnaseer):

ok, you agree the following is true?\[(x+6)^2=x^2+12x+36\]

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

good, so next we did this:\[(x+6)^2-18=x^2+12x+36-18=x^2+12x+18\]make sense so far?

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

right, now your original equation is just the right-hand-side of what I just wrote

OpenStudy (asnaseer):

so your original:\[x^2+12x+18=0\]can be rewritten as:\[(x+6)^2-18=0\]do you see?

OpenStudy (anonymous):

yeah

OpenStudy (asnaseer):

good, so next we just add 18 to both sides to get:\[(x+6)^2=18\]

OpenStudy (anonymous):

oh

OpenStudy (asnaseer):

and you know how to solve from there

OpenStudy (anonymous):

tap re-do the problem and you will see it. Dont just follow him along. you didnt follow how he got to that part...

OpenStudy (asnaseer):

I can try and explain it more simply by working through another example if you want?

OpenStudy (anonymous):

yes please…do you want another from my homework or were you going to make it up?

OpenStudy (asnaseer):

either - which do you prefer?

OpenStudy (anonymous):

homework. 7t^2+28t+56=0

OpenStudy (asnaseer):

ok, so if you recall the steps I listed above, these were: 1) divide by the coefficient of x^2 first so here we have:\[7t^2+28t+56=0\]so 1st we divide by 7 to get:\[t^2+4t+8=0\]step 2) was to half the coefficient of the x-term to form the square, so here we get:\[(t+2)^2=t^2+4t+4\]but we want:\[t^2+4t+8\]so we need to add on another 4 to get:\[(t+2)^2+4=t^2+4t+8\]so we can now use this in the original to get:\[(t+2)^2+4=0\]giving:\[(t+2)^2=-4\]

OpenStudy (asnaseer):

do the steps make sense?

OpenStudy (anonymous):

yes. thanks so much.

OpenStudy (asnaseer):

no problem - I'm glad we got there in the end :-)

OpenStudy (asnaseer):

the basic step involved here is, when you have something like:\[x^2+bx+c=0\]to recognise that:\[(x+p)^2=x^2+2px+p^2\]and use this

OpenStudy (asnaseer):

here 2p must equal b -- which is why we have to half the coefficient of the x-term when forming the square

OpenStudy (anonymous):

ok. have to go to dance. thanks again!!

OpenStudy (asnaseer):

ok - enjoy your dancing...